OpenStudy (anonymous):

how do you factor the equation x^2+5x+6=0

sam (.sam.):

In this problem 3*2=6 and 3+2=5, so insert 3 as the right hand term of one factor and 2 as the right-hand term of the other factor. (x+3)(x+2)=0 Set each of the factors of the left-hand side of the equation equal to 0. x+3=0,x+2=0 The complete solution is x=-3,-2

OpenStudy (anonymous):

x2 + 5x + 6 = (x + 2)(x + 3)

OpenStudy (anonymous):

(x + 2)(x + 3) = 0 so : (x + 2)=0 , (x + 3) = 0 So: x=-2 , x=-3